OpenStudy (anonymous):

The radioactive substance uranium-240 has a half-life of 14 hours. The amount At of a sample of uranium-240 remaining (in grams) after t hours is given by the following exponential function.

OpenStudy (anonymous):

\[A(t) = 4700\left(\begin{matrix}1 \\ 2\end{matrix}\right)^{\frac{ t }{ 14 }}\]

OpenStudy (anonymous):

Find the initial amount in the sample and the amount remaining after 60 hours. Round your answers to the nearest gram as necessary.

OpenStudy (anonymous):

Is that a I should know this jeez it's understandable why I need help on this jeez lol I'm a senior though

OpenStudy (anonymous):

or*

OpenStudy (anonymous):

yeahh lol I don't think mathway can help much until I get it started and figure out how to narrow it down to a better equation haha but thank you =)

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Is the formula \(A(t) = 4700\left(\begin{matrix}1 \\ 2\end{matrix}\right)^{\frac{ t }{ 14 }}\) or is it \(\large A(t) = 4700(\frac{1}{2})^{\frac{t}{14}}\)

OpenStudy (anonymous):

the second one sorry

OpenStudy (mathmate):

Use a calculator to find A(0) (initial amount, put t=0), and A(60), i.e. put t=60 hours.

OpenStudy (anonymous):

I'm a little confused I don't understand what I'm calculating and what it is going to answer out of the two questions =o @mathmate

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

I'm not good at half-life.. sorry!

OpenStudy (anonymous):

lol I've never even heard of it =/ but it's ok thank you anyways =)

OpenStudy (mathmate):

@Cassandra_Lea_96 To give an example, the quantity remaining after t=2 hours is A(2)=4700(1/2)^(t/14)=4700(1/2)^(2/14)=4257 approximately (you'll need a calculator to calculate that). So the answers will be similarly calculated for A(0) and A(60).

OpenStudy (anonymous):

@mathmate can you help me out there please

OpenStudy (mathmate):

@Cassandra_Lea_96 Have you done functions before? Like, if f(x)=x^2+4, then f(0)=0^2+4=4, f(2)=2^2+4=8, and so on.

OpenStudy (anonymous):

I've done functions before but this is just completely confusing me aha is the initial amount 4700? and then that times 60 is the ammount after 60 days or

OpenStudy (anonymous):

ok so is it 240.97? it won't let me type in a fraction or anything so this is it in decimal form

OpenStudy (mathmate):

Yes, you've got both answers correct. Decimal form is good. Well done!